# Post-processing improves accuracy of Artificial Intelligence weather forecasts

### v1

@@ Abstract
Artificial Intelligence (AI) weather models are now reaching operational-grade performance for some variables, but like traditional Numerical Weather Prediction (NWP) models, they exhibit systematic biases and reliability issues. We test the application of the Bureau of Meteorology’s existing statistical post-processing system, IMPROVER, to ECMWF’s deterministic Artificial Intelligence Forecasting System (AIFS), and compare results against post-processed outputs from the ECMWF HRES and ENS models. Without any modification to configuration or processing workflows, post-processing yields comparable accuracy improvements for AIFS as for traditional NWP forecasts, in both expected value and probabilistic outputs. We show that blending AIFS with NWP models improves overall forecast skill, even when AIFS alone is not the most accurate component. These findings show that statistical post-processing methods developed for NWP are directly applicable to AI models, enabling national meteorological centres to incorporate AI forecasts into existing workflows in a low-risk, incremental fashion.
@@
This Work has been submitted to Artificial Intelligence for the Earth Systems. Copyright in this Work may be transferred without further notice.
@@ 1 Introduction
Artificial Intelligence weather models have rapidly improved in the last few years (de Burgh-Day and Leeuwenburg, 2023). Deterministic AI models now typically achieve RMSE accuracy as good or better than traditional Numerical Weather Prediction (NWP) models for medium-range forecasting on most variables assessed, while being many times faster to run. In this study, we focus on the European Centre for Medium-Range Weather Forecasting’s (ECMWF) deterministic Artificial Intelligence Forecasting System (AIFS) model (Lang et al., 2024a), which is operationally supported as of 25 Febrary 2025 (ECMWF, 2025b). AIFS is a graph transformer neural network. Graph-based architectures have been hugely successful in weather modelling. Deepmind’s GraphCast model was found to have better RMSE skill score than ECMWF’s HRES deterministic model in around 90% of cases (Lam et al., 2023). Like GraphCast, AIFS generally outperforms HRES (Lang et al., 2024a). Earlier work by Keisler (2022) used a much smaller graph-based network and achieved results better than the Global Forecasting System (GFS, produced by the US National Oceanic and Atmospheric Administration) but not as good as the ECMWF’s HRES model.
As is the case for traditional NWP models, the resolution of AI weather models is limited by computational constraints.The resolution of AIFS is approximately 0.25°0.25°0.25  ° (Lang et al., 2024a), significantly coarser than the 0.1°0.1°0.1  ° resolution of ECMWF HRES (ECMWF, 2025a). Also, as discussed in section 4, AI models, and graph-based models in particular, tend to produce gridded outputs that are spatially very smooth. For these reasons, AI models can have difficulty modelling small-scale phenomena. Furthermore, the graph-based models described above are all auto-regressive: the first timestep is calculated by running the model on the current analysis, and then subsequent timesteps take as input the model output of the previous timestep. Therefore, it is expected that model errors can accumulate over time. Because of these limitations, we hypothesise that post-processing methods developed for NWP can be applied to AI forecasts without modification and yield comparable results, and that incorporating AI models into forecast blends will improve both deterministic and probabilistic forecast skill.
@@
Bülte et al. (2025) evaluate two post-processing methods for producing probabilistic forecasts from Pangu-Weather, and compare the results with the ECMWF raw ensemble. In contrast to the work of Bremnes et al. (2024), the models are trained and evaluated on gridded analyses rather than site data. The first method is the EasyUQ technique developed by Walz et al. (2024) and based on isotonic distributional regression of Henzi et al. (2021). This is a non-parametric non-AI calibration method somewhat similar to the reliability calibration used in the present work. The second method is distributional regression networks, an AI approach where post-processed forecasts follow parametric distributions whose parameters are predicted by a neural network based on the inputs. While the latter method allows incorporating additional predictors, it is found not to offer significant advantages over the simpler non-AI method. Both methods allow the post-processed deterministic forecast to achieve similar or better accuracy than the raw ECMWF ensemble, at least for the first few days of the forecast period. A much more sophisticated method is demonstrated by Ge et al. (2022), who use a deep neural network to bias correct and downscale gridded predictions to the point scale. Since the network has billions of parameters, training is computationally intensive and requires a long history of training data.
Here we apply the IMPROVER (Roberts et al., 2023) post-processing sytem to post-process the AIFS forecast outputs, and apply the same methods to ECMWF’s deterministic HRES and ensemble (ENS) models to provide a comparison. We demonstrate that although IMPROVER is developed for post-processing traditional NWP forecasts, it is also effective for AIFS with no changes to the configuration parameters or processing workflows. We also show that adding AIFS to a blend of models with the two tradtional NWP forecasts improves the accuracy of the blend, even in situations where AIFS is less accurate than one of the other models. One of the key features of IMPROVER is that it facilitates producing probabilistic forecasts from both deterministic and ensemble NWP inputs. The probabilistic forecasts produced by AIFS are of comparable quality to those produced by ECMWF HRES and including AIFS in the blend improves the quality of the probabilistic forecast.
@@ 2 Data
The study uses data spanning 1 March 2024 to 23 July 2024 (with a few days of missing data, varying by parameter). The date range was limited by computational resources and the availability of archived pre-processed data. In future work, it would be interesting to consider a longer date range including the Summer period. AIFS data was downloaded from ECMWF’s Meteorological Archival and Retrieval System (MARS). Since calibration uses a rolling period of the previous 30 days of forecasts, the outputs are not reliable for the first month, and we analyse the results only from 1 April onwards, amounting to approximately 16 weeks of data. We consider forecasts from the 12Z basetime for each day in the period. We selected this basetime since, of the Bureau’s two daily operational forecasts, this one receives the most attention. We post-process 3 surface-level variables: temperature, dew point temperature, and wind speed at 10m. We evaluate both the expected value and probabilistic output forecasts.
Post-processed forecasts are produced for lead times up to 240 hours at 1-hour intervals. The raw AIFS forecast has lead times at 6-hourly intervals, while the ENS and HRES forecasts have 3-hourly frequency up to 150 hours, and 6-hourly thereafter. The resolutions of the NWP forecasts ENS and HRES are 0.2°0.2°0.2  ° and 0.1°0.1°0.1  °, respectively. [^1] ENS is produced by ECMWF at 0.1°0.1°0.1  ° resolution; the Bureau receives an upscaled version. The AIFS forecast is produced on the N320 reduced Gaussian grid (see Lang et al. (2024a)). Our bias correction and calibration use the gridded analysis produced by Mesoscale Surface Analysis System (MSAS) (Glowacki et al., 2012), which has a resolution of 2.5 arc minutes (approximately 4 km on the Australian domain).
Hourly observation data was extracted from the Bureau of Meteorology’s Jive database (Loveday et al., 2024), which contains quality-controlled observations from the Bureau’s network of Automatic Weather Stations. In total, 569 weather stations are used in the analysis. Figure 1 shows the location of these stations.
@@ 3 Methods
AIFS data was pre-processed to a format expected by IMPROVER. As mentioned above, AIFS is produced on the N320 reduced Gaussian grid. The IMPROVER software cannot natively handle this grid, so as part of pre-processing we transform it to a latitude/longitude grid of resolution 0.25°0.25°0.25  °. IMPROVER is already in use operationally at the Bureau of Meteorology for HRES, ENS, and MSAS, so for these models pre-processed data was retrieved from an operational archive.
IMPROVER produces probabilistic forecasts by predicting the probabilities of exceedence at a set of thresholds. Expected value forecasts were calculated from the probabilistic output by linearly interpolating these probabilities and integrating the resulting distribution.
The Bureau’s IMPROVER processing workflow is summarised in Section 3.1 and described in detail in (Owen et al., 2024) and (Owen et al., 2025). Forecast bias and probability calibration are corrected using a rolling history of 30 days of data. The ground truth for these corrections is the gridded MSAS analysis mentioned above. Probability calibration is implemented using the reliability calibration algorithm described by Flowerdew (2014).
After running the IMPROVER hindcast with AIFS, the output was extracted at the observation sites (Figure 1), and expected values were calculated from the thresholded forecast. For temperature, the expected value (but not the thresholded site forecast) was adjusted using the standard adiabatic lapse rate of -0.0098 °C/m to account for the difference between the grid point average altitude and the site altitude.
@@ 3.1 IMPROVER processing
Below we describe the main post-processing steps implemented by our IMPROVER operational configuration. More detail can be found in (Owen et al., 2024) and (Owen et al., 2025).
@@
    Regridding to 2400m Albers equal area grid. (Note that this is in addition to the initial regridding of AIFS from the N320 grid as mentioned above.)
@@
    Thresholding (ensemble forecast) or fuzzy thresholding (deterministic forecast) to produce a probabilistic forecast. IMPROVER represents probabilistic forecasts as a set of probabilities and their corresponding probabilities of exceedence. For an ensemble forecast, thresholding simply means calculating the proportion of ensemble members that exceed the threshold. For a deterministic forecast, simple thresholding would yield a probability of either 0 or 1, which is not useful for reliability calibration. Instead, we use the fuzzy thresholding approach, where the predicted value is mapped to a number between 0 and 1 depending how far it is from the threshold. See Owen et al. (2025) for more details on this technique.
@@
    Neighbourhood processing with 2500 m radius. This smooths the probabilistic forecast by convolving it with a constant-valued square filter.
@@
    Recursive filtering. An additional smoothing operation, similar to an exponential weighted average filter.
@@
    Reliability calibration. Probabilities are adjusted using a piecewise-linear correction function calculated as follows. For each threshold, the previous 30 days’ history of forecasts are binned into 7 probability bins, and for each bin, the average forecast and observed probabilities of exceeding the threshold are calculated. The forecast and observed probabilities define the $x$ and $y$ coordinates, respectively, of the knot points of the piecewise linear correction function. The historical forecasts are archived after the thresholding step. Reliability calibration is calculated separately for each lead time, using the aggregate of all grid points. As is the case for bias correction, the MSAS analysis is used as ground truth.
@@ 3.2 Blending weights
We consider two blended forecasts: the first uses only the two NWP models, ENS and HRES, and the second uses all three models. In each case, the blending weights are chosen as follows. The data is split in half by valid time. Weights are fitted separately on each half of the data, and then applied to create the blended forecast for the other half of the data. Each set of weights is calculated as follows. First, we calculate the optimal weights at each lead time to minimise the mean squared error. However, these weights vary greatly from hour to hour, partly due to diurnal patterns in the model biases, and partly due to random noise. To reduce overfitting and improve generalisation, we calculate a second set of smoothed weights. The smoothed weight for each model is a piecewise-linear function of lead time with 11 knots at lead times 0, 24, 48, … 240. The function values at the knot points are chosen to minimise the MSE between the piecewise linear function and the individual hourly optimal weights, then normalised to sum to 1 over all models at each knot point (which guarantees that the interpolated values between knot points also sum to 1).
@@
Figure 2: Blending weights for all-model blend (top row) and NWP-model blend (ENS and HRES only; bottom row). The dotted line shows the optimal weights per lead hour. The solid line shows a smoothed version where a piecewise-linear function is fitted to the optimal weights.

### v2

# Statistical post-processing yields accurate probabilistic forecasts from Artificial Intelligence weather models

@@ Abstract
Artificial Intelligence (AI) weather models are now reaching operational-grade performance for some variables, but like traditional Numerical Weather Prediction (NWP) models, they exhibit systematic biases and reliability issues. We test the application of the Bureau of Meteorology’s existing statistical post-processing system, IMPROVER, to ECMWF’s deterministic Artificial Intelligence Forecasting System (AIFS), and compare results against post-processed outputs from the ECMWF HRES and ENS models. Without any modification to processing workflows, post-processing yields comparable accuracy improvements for AIFS as for traditional NWP forecasts, in both expected value and probabilistic outputs. We show that blending AIFS with NWP models improves overall forecast skill, even when AIFS alone is not the most accurate component. These findings show that statistical post-processing methods developed for NWP are directly applicable to AI models, enabling national meteorological centres to incorporate AI forecasts into existing workflows in a low-risk, incremental fashion.
@@
This Work has been accepted by Artificial Intelligence for the Earth Systems. The AMS does not guarantee that the copy provided here is an accurate copy of the Version of Record (VoR).
@@ 1 Introduction
Artificial Intelligence weather models have rapidly improved in the last few years (de Burgh-Day and Leeuwenburg, 2023). Deterministic AI models now typically achieve RMSE accuracy as good or better than traditional Numerical Weather Prediction (NWP) models for medium-range forecasting on most variables assessed, while being many times faster to run. In this study, we focus on the European Centre for Medium-Range Weather Forecasting’s (ECMWF) deterministic Artificial Intelligence Forecasting System (AIFS) model (Lang et al., 2024a), which is operationally supported as of 25 February 2025 (ECMWF, 2025b). AIFS is a graph transformer neural network. Graph-based architectures have been hugely successful in weather forecasting. Deepmind’s GraphCast model was found to have better RMSE skill score than ECMWF’s HRES deterministic model in around 90% of cases (Lam et al., 2023). Like GraphCast, AIFS generally outperforms HRES (Lang et al., 2024a). Earlier work by Keisler (2022) used a much smaller graph-based network and achieved results better than the Global Forecasting System (GFS, produced by the US National Oceanic and Atmospheric Administration) but not as good as the ECMWF’s HRES model.
As is the case for traditional NWP models, the resolution of AI weather models is limited by computational constraints. The resolution of AIFS is approximately 0.25° (Lang et al., 2024a), significantly coarser than the 0.1° resolution of ECMWF HRES (ECMWF, 2025a). Furthermore, the graph-based models described above are all auto-regressive: the first timestep is calculated by running the model on the current analysis, and then subsequent timesteps take as input the model output of the previous timestep. Therefore, model errors can accumulate over time. Thus, as with NWP models, a key objective for post-processing AI models is to downscale and bias-correct the forecast. Additionally, operational post-processing systems now commonly produce probabilistic as well as deterministic outputs, and combine various input models into a blended output forecast. We hypothesise that existing post-processing systems, developed for NWP models, can be used without modification to fulfill all these objectives for AI models.
@@
Bülte et al. (2025) evaluate two post-processing methods for producing probabilistic forecasts from Pangu-Weather, and compare the results with the ECMWF raw ensemble. In contrast to the work of Bremnes et al. (2024), the models are trained and evaluated on gridded analyses rather than site data. The first method evaluated by Bülte et al. (2025) is the EasyUQ technique developed by Walz et al. (2024) and based on isotonic distributional regression of Henzi et al. (2021). This is a non-parametric calibration method somewhat similar to the reliability calibration used in the present work (and described in more detail in Section 33.1). The second method is distributional regression networks, an approach where post-processed forecasts follow parametric distributions whose parameters are predicted by a neural network based on the inputs. While the latter method allows incorporating additional predictors, it is found not to offer significant advantages over the simpler method. Both methods allow the post-processed deterministic forecast to achieve similar or better accuracy than the raw ECMWF ensemble, at least for the first few days of the forecast period. A much more sophisticated method is demonstrated by Ge et al. (2022), who use a deep neural network to bias correct and downscale gridded predictions to the point scale. Since the network has billions of parameters, training is computationally intensive and requires a long history of training data.
Here we apply the IMPROVER (Roberts et al., 2023) post-processing system to post-process the AIFS forecast outputs, and apply the same methods to ECMWF’s deterministic HRES and ensemble (ENS) models to provide a comparison. We demonstrate that although IMPROVER is developed for post-processing traditional NWP forecasts, it is also effective for AIFS with no changes to the configuration parameters or processing workflows. We also show that adding AIFS to a blend of models with the two traditional NWP forecasts improves the accuracy of the blend, even in situations where AIFS is less accurate than one of the other models. One of the key features of IMPROVER is that it facilitates producing probabilistic forecasts from both deterministic and ensemble NWP inputs. The probabilistic forecasts produced by AIFS are of comparable quality to those produced by ECMWF HRES and including AIFS in the blend improves the quality of the probabilistic forecast.
@@ 2 Data
The study uses data spanning 1 March 2024 to 23 July 2024 (with a few days of missing data, varying by parameter). The date range was limited by computational resources and the availability of archived pre-processed data. In future work, it would be interesting to consider a longer date range including the Southern hemisphere summer period. AIFS data was downloaded from ECMWF’s Meteorological Archival and Retrieval System (MARS). Since calibration uses a rolling period of the previous 30 days of forecasts, the outputs are less accurate for the first month, and we analyse the results only from 1 April onwards, amounting to approximately 16 weeks of data. We consider forecasts from the 12Z basetime for each day in the period. [^1] Basetime refers to the time that the model’s initial conditions represent. We selected this basetime since, of the Bureau’s two daily operational forecasts, this one receives the most attention. We post-process 3 surface-level variables: temperature, dew point temperature, and wind speed at 10m. We evaluate both the expected value and probabilistic output forecasts.
Post-processed forecasts are produced for lead times up to 240 hours at 1-hour intervals. The raw AIFS forecast has lead times at 6-hourly intervals, while the ENS and HRES forecasts have 3-hourly frequency up to 150 hours, and 6-hourly thereafter. As described in Section 3, the input forecasts are interpolated to the output frequency. The resolutions of the NWP forecasts ENS and HRES are 0.2° and 0.1°, respectively. [^2] ENS is produced by ECMWF at 0.1° resolution; the Bureau receives an upscaled version. The AIFS forecast is produced on the N320 reduced Gaussian grid (see Lang et al. (2024a)). Our bias correction and calibration use the gridded analysis of the Australian domain produced by the Mesoscale Surface Analysis System (MSAS) (Glowacki et al., 2012), which has a resolution of 2.5 arc minutes (approximately 4 km for the region modelled).
The ground truth for our verification is hourly observation data extracted from the Bureau of Meteorology’s Jive database (Loveday et al., 2024), which contains quality-controlled observations from the Bureau’s network of Automatic Weather Stations. In total, 569 weather stations are used in the analysis. Figure 1 shows the location of these stations. Observation data is also used for optimising the blending weights as discussed in Section 33.2.
@@ 3 Methods
The full analysis workflow is summarised in Figure 2.
InputRegrid to lat/lon(AIFS only)IMPROVER processingGriddedprobabilitiesExtract sitesProbabilitiesat sitesCalculateexpected valuesExpected valuesat sitesBlend probabilitieswith other forecastsBlend expected valueswith other forecastsInterpolatetime stepsRegridLapse rate correction(temperature only)Bias correction(Fuzzy) thresholding(transform forecastto probabilities)Spatial smoothingReliablity calibration
Figure 2: The processing workflow. The flowchart on the left shows the main steps of the analysis. These steps are duplicated for each input model (ENS, HRES, and AIFS). Inputs and outputs of selected steps are shown in bold. The IMPROVER processing is expanded in more detail on the right.
AIFS data was pre-processed to a format expected by the IMPROVER software. As mentioned above, AIFS is produced on the N320 reduced Gaussian grid. The IMPROVER software cannot natively handle this grid, so as part of pre-processing we transform it to a latitude/longitude grid of resolution 0.25°. IMPROVER is already in use operationally at the Bureau of Meteorology for HRES, ENS, and MSAS, so for these models pre-processed data was retrieved from an operational archive. [^3] This pre-processing takes place prior to the IMPROVER workflow described below and is very minimal. It primarily consists of standardising file formats (including converting from GRIB to NetCDF format, standardising variable names, and extracting the Australian region). The inputs and outputs for IMPROVER cover the Australian region only; for the global models HRES, ENS, and AIFS, the Australian domain is extracted in pre-processing that takes place before IMPROVER.
The IMPROVER software produces probabilistic forecasts by predicting the probabilities of exceedence at a set of thresholds. The thresholds are chosen so that they cover the range of forecast values one could reasonably expect for the Australian climate. We use quite a large number of thresholds so that the probability distributions can be modelled realistically: there are 61 thresholds for temperature, 47 for dewpoint, and 49 for windspeed. Expected value forecasts are calculated from the probabilistic output by linearly interpolating these probabilities and integrating the resulting distribution.
The Bureau’s IMPROVER processing workflow is summarised in Section 3.1 and described in detail in Owen et al. (2024) and Owen et al. (2025). Forecast bias and probability calibration are corrected using a rolling history of 30 days of data. The ground truth for these corrections is the gridded MSAS analysis mentioned above. Probability calibration is implemented using the reliability calibration algorithm described by Flowerdew (2014).
After running the IMPROVER processing with AIFS, the output was extracted at the observation sites (Figure 1), using the nearest-neighbour method. [^4] The IMPROVER software was also used to extract the site forecasts, but this was achieved with a post-processing script separate from the main IMPROVER processing workflow. This is because the Bureau’s operational IMPROVER workflow scheduler, which we adapted for this analysis, produces only gridded forecasts. Expected values were calculated from the thresholded forecast. [^5] Again, this was done using the IMPROVER software, but as a post-processing step separate from the main processing of gridded forecasts. This allowed us to calculate expected values only for site data, rather than gridded data, greatly speeding up the analysis. For temperature, the expected value (but not the thresholded site forecast) was adjusted using the dry adiabatic lapse rate of -0.0098 °C/m to account for the difference between the grid point average altitude and the site altitude.
@@ 3.1 IMPROVER processing
IMPROVER produces gridded forecasts over the Australian domain. Below we describe the main post-processing steps implemented by our IMPROVER operational configuration. More detail can be found in Owen et al. (2024) and Owen et al. (2025).
@@
    Regridding to 2400m Albers equal area grid using bilinear interpolation. (Note that this is in addition to the initial regridding of AIFS from the N320 grid as mentioned above.)
@@
    Thresholding (ensemble forecast) or fuzzy thresholding (deterministic forecast) to produce a probabilistic forecast. IMPROVER represents probabilistic forecasts as a set of thresholds and their corresponding probabilities of exceedence. For an ensemble forecast, thresholding simply means calculating the proportion of ensemble members that exceed the threshold. For a deterministic forecast, simple thresholding would yield a probability of either 0 or 1, which is not useful for reliability calibration. Instead, we use the fuzzy thresholding approach, where the predicted value is mapped to a number between 0 and 1 depending on how far it is from the threshold. Specifically, a parameter $r$ is chosen, representing the distance from the threshold within which we want to measure variation, and then for a deterministic forecast $f$ and a threshold $t$, the output $f^{*}$ of fuzzy thresholding is the number between 0 and 1 defined as follows
    $f^{*}=\begin{cases}0,&x<t-r,\\ (f-t+r)/(2r),&t-r\leq f\leq t+r,\\ 1,&x>t+r\end{cases}$
    See Owen et al. (2025) for more details on this technique.
@@
    Neighbourhood processing with neighbourhoods of size $3\times 3$ grid cells. This smooths the probabilistic forecast by convolving it with a constant-valued square filter. See Roberts (2003) for more detail.
@@
    Recursive filtering. An additional spatial smoothing operation, similar to an exponential weighted average filter. See Roberts (2003) for more detail.
@@
    Reliability calibration. Probabilities are adjusted using the reliability calibration method of Flowerdew (2014). A piecewise-linear correction function is calculated as follows. For each threshold, the previous 30 days’ history of forecasts are binned into 7 probability bins, and for each bin, the average forecast and observed probabilities of exceeding the threshold are calculated. The forecast and observed probabilities define the $x$ and $y$ coordinates, respectively, of the knot points of the piecewise linear correction function. The number of bins was chosen to give enough flexibility to model the calibration curve while avoiding overfitting (due to small sample size in some bins) and keeping the computation time feasible. The first bin contains only the value zero, and the remaining bins divide the interval $(0,1]$ into equally-sized intervals. Our binning arrangement is similar to that used in Rust et al. (2023), who also use 7 bins in total, including two single-value bins at 0 and 1. The historical forecasts are archived after the thresholding step. Reliability calibration is calculated separately for each lead time, using the aggregate of all grid points. Thus (unlike the bias correction) the reliability calibration can address only global calibration problems, not those specific to particular locations or regions. Calculating separate reliability calibrations for each grid point would not be computationally feasible, and would also risk overfitting, given the short calibration period. As is the case for bias correction, the MSAS analysis is used as ground truth.
@@ 3.2 Blending weights
Forecast blending is the practice of combining outputs of multiple forecasts. In our case the blends are simple weighted averages of either the probability distribution outputs or the expected value outputs of post-processing. In order to investigate the benefits of adding AI models to a blend of NWP models, we compare two blended forecasts: the first uses only the two NWP models, ENS and HRES, and the second uses all three models. To supplement this analysis, we also consider the two other two-model blends, namely AIFS with HRES, and AIFS with ENS. In each case, the blending weights are chosen as follows. The data is split in half by valid time. [^6] Valid time refers to the time the forecast is for. Weights are fitted separately on each half of the data, and then applied to create the blended forecast for the other half of the data (this approach is often called 2-fold cross-validation). In this way, we avoid fitting the weights on the same data we use to measure model performance. In general, the performance of a weighting model is expected to be slightly better when applied to the same dataset used to optimise the weights, compared to an unseen dataset, because the weights can fit to random noise in the training dataset. Thus our cross-validation approach is intended to obtain a realistic indication of the model’s performance in an operational setup, where the weights must be fitted on historical data, and then applied on previously-unseen data.
The ground truth for fitting the weights is the observation data at sites, and data is pooled so that weights are shared across all sites. To produce weights that vary smoothly over lead times, we use the following process. First, we calculate the optimal weights at each lead time to minimise the mean squared error. However, these weights vary greatly from hour to hour, partly due to diurnal patterns in the model biases, and partly due to random noise. To reduce overfitting and improve generalisation, we calculate a second set of smoothed weights. The smoothed weight for each model is a piecewise-linear function of lead time with 11 knots at lead times 0, 24, 48, … 240. The function values at the knot points are chosen to minimise the MSE between the piecewise linear function and the individual hourly optimal weights, then normalised to sum to 1 over all models at each knot point (which guarantees that the interpolated values between knot points also sum to 1).
@@
Figure 3: Blending weights for the all-model blend (top row) and NWP-model blend (ENS and HRES only; bottom row). The dotted line shows the optimal weights per lead hour. The solid line shows a smoothed version where a piecewise-linear function is fitted to the optimal weights. These weights are calculated on the first half of the dataset, and are used to calculate the blended output for the second half.